OpenStudy (anonymous):

2 1/2 of milk makes 5 servings.How many cup of milk are needed to make 8 servings?

OpenStudy (lucaz):

notice 2 1/2 is 5/2, from 5/2 to 5 you multiply by 2, from x to 8 you also multiply by 2, so x = 8/2 = 4

OpenStudy (anonymous):

\[2.5\rightarrow5\] \[x \rightarrow 8\] \[5x \rightarrow 2.5*8\] x\[x \rightarrow 5*2.5*8 = 4 \] so 4 cup milk needed